Polyethylene is a synthetic polymer or plastic with many uses. 1.44 g of a polyethylene sample was dissolved in enough benzene to make 100. mL of solution, and the osmotic pressure was found to be 1.86 torr at 25oC. What is the approximate molar mass of the polyethylene?
Answer #1

PI = MRT

PI = osmotic pressure

M = concnetration in mol per liter

M = mol/L

L = 100 ml = 0.1 L

R= 0.082 ideal gas constant

T = 25 °C = 298 K

then

PI = mol/L * R

mol = PI*L/(RT)

PI = 1.86 torr = 1.86/760 = 0.00244736842 atm

mol = 0.00244736842/0.1/(0.082*298) = 0.00100154215 mol of plastic

MW = mass/mol

MW = 1.44/0.00100154215 = 1437.782 g /mol of polyethylene